# A Subject-Agnostic and Versatile Model for fMRI-to-Text Decoding

### v1

@@ Abstract
Decoding functional magnetic resonance imaging (fMRI) signals into text has been a key challenge in the neuroscience community, with the potential to advance brain-computer interfaces and uncover deeper insights into brain mechanisms. However, existing approaches often struggle with suboptimal predictive performance, limited task variety, and poor generalization across subjects. In response to this, we propose MindLLM, a model designed for subject-agnostic and versatile fMRI-to-text decoding. MindLLM consists of an fMRI encoder and an off-the-shelf LLM. The fMRI encoder employs a neuroscience-informed attention mechanism, which is capable of accommodating subjects with varying input shapes and thus achieves high-performance subject-agnostic decoding. Moreover, we introduce Brain Instruction Tuning (BIT), a novel approach that enhances the model’s ability to capture diverse semantic representations from fMRI signals, facilitating more versatile decoding. We evaluate MindLLM on comprehensive fMRI-to-text benchmarks. Results demonstrate that our model outperforms the baselines, improving downstream tasks by , unseen subject generalization by , and novel task adaptation by . Furthermore, the attention patterns in MindLLM provide interpretable insights into its decision-making process.
@@ 1 Introduction
Another critical challenge lies in designing a unified and subject-agnostic architecture. Current methods of brain multimodal decoding mostly rely on a preprocessing step: selecting responsive voxels by comparing the task-based fMRI to the resting-state fMRI. The selection typically results in higher performance compared to using whole-brain data. However, the varying number and irregular spatial distribution of selected voxels across subjects pose significant challenges for developing a unified architecture. To this end, recent studies (Wang et al., 2024a, b) have proposed pooling or sampling voxels to standardize input dimensions. However, as illustrated in Figure 3, these methods still suffer from the loss of spatial information and uneven representations of certain areas, ultimately comprising performance.
@@ 2 Related Works
This line of research mostly focuses on decoding perceived visual stimuli into natural language from fMRI signals. MindGPT (Chen et al., 2023a), UniBrain(Mai & Zhang, 2023) and BrainCap (Ferrante et al., 2023) employ an fMRI encoder guided by CLIP (Radford et al., 2021) and use a language model (Radford et al., 2019; Wang et al., 2022) to decode natural language from the encoded representations. BrainChat (Huang, 2024) utilizes multiple pretraining strategies (Devlin, 2018; He et al., 2022; Yu et al., 2022) to align fMRI with image and text embeddings. These methods fall short in performance and versatility. UMBRAE (Xia et al., 2024) proposes to learn a mapping from fMRI to stimulus images, which later serves as a proxy input for an off-the-shelf visual language model (VLM). Although they achieve performance improvements, the strategy prevents the model from performing tasks that are not directly related to the stimulus images (e.g. answering memory-related questions). In contrast, our end-to-end Brain Instruction Tuning (BIT) ensures seamless and versatile fMRI-to-text decoding, offering the potential to tackle tasks beyond vision-related ones.
@@ 3.1 Method Overview
As mentioned before, currently most models for fMRI decoding can not handle varying input shapes and are not subject-agnostic, with only a few exceptions (Mai & Zhang, 2023). However, these exceptions still suffer from information loss and uneven representations of certain brain areas. To this end, we propose a novel neuroscience-informed attention mechanism to accommodate varying voxel numbers across subjects, enabling a subject-agnostic encoding strategy. Below we talk about the design of queries , keys and values in the attention layer. For values, we directly use the fMRI signal of each voxel, which means . Making each voxel a value token maximally prevents information loss compared to pooling- (Wang et al., 2024a) or sampling-based (Mai & Zhang, 2023) methods. The queries are randomly initialized and learnable. We expect each query to represent a certain pattern of the brain (refer to visualizations in Section 4.7). The design of keys will be discussed below.
Exclude fMRI values from keys The vanilla cross attention (Zhu et al., 2020; Vaswani, 2017) derives both keys and values from the same input source. However, we found this would lead to poor performance in fMRI. We argue the reason: different from images or text, which are usually considered translation-invariant, the positions of voxels carry specific brain functional information, as voxels in different areas are associated with distinct brain functions. Consequently, a voxel’s position alone can theoretically serve as effective keys for attention weight computation. Including fMRI values into keys, however, introduces additional noise instead of valuable information, thus resulting in poorer performance. Moreover, since brain regions tend to serve similar functions across individuals, decoupling voxel positions from fMRI signals can facilitate the sharing of priors across subjects, potentially improving generalization to unseen subjects.
@@
Combining Multiple Parcellations It is crucial to choose an appropriate brain region parcellation. Previous region-based methods (Qiu et al., 2023; Li et al., 2021; Kan et al., 2022) can usually only choose one arbitrarily. In contrast, our model design allows us to combine multiple parcellations by concatenating their respective region encodings to the key embeddings. In conclusion, the final key embeddings are the concatenation by the positional encoding and multiple region encodings,
@@ 4.1 Settings
Downstream Datasets The downstream dataset will be discussed within each experiment section. See examples and a short description for all dataset we will use in Appendex A. Implementation details could be found in Appendix B.
@@ 4.3 Versatile Decoding
Baselines Besides baselines that could be adapted to this experiment from the previous one, we further consider the following subject-agnostic models as baselines. 1) MindBridge (Wang et al., 2024a) flatten the voxels and adaptively adjust the padding and stride to pool the voxels into a fixed dimension. The original implementation of MindBridge has subject-specific parameters. However, since those parameters are of the same size, we make them shared across subjects and thus make the model subject-agnostic. 2) UniBrain (Wang et al., 2024b) samples voxels into a fixed number of groups and employs a transformer where groups are treated as tokens. This UniBrain is unrelated to the UniBrain in the previous section; they just share the same name.
@@
The results are shown in Table 2. Our model outperforms baselines, with an average improvement of . Further, by comparing instruction tuning and from-scratch models, we find that instruction tuning has a significant positive effect, with an average improvement of . The results remain stable across different random seeds; for instance, according to our observations, the BLEU-1 score for paragraph captioning exhibits a maximum of variance.
@@ 4.4 Unseen Subject Generalization
Our neuroscience-informed, subject-agnostic design enhances generalization to novel subjects, a crucial factor in real-world applications where training a model for each individual is impractical. To evaluate it, we perform instruction tuning on out of the subjects in the natural scene dataset (Allen et al., 2022), and evaluate generalization on the held-out subject. Table 3 shows our model outperforms two other subject-agnostic baselines in most cases, with an average improvement of compared to the second-best model.
@@ 4.6 Ablation Study
We conduct ablation studies on the design of key embeddings in the neuroscience-informed attention module in Figure 7. The results strongly validate our design. The vanilla cross attention (Pos Enc.+fMRI) leads to poor performance while removing fMRI values from the key embeddings (Pos Enc.) yields a significant improvement. Replacing positional encoding with region encodings (Reg. Enc.) accelerates convergence in the early stages since it is grounded by neuroscientific principles. However, it is eventually outperformed by Pos Enc. due to the lack of finer-grained information. Combining the positional encoding and region encodings (Pos Enc.+Reg Enc.) achieves the best model design. In addition, replacing positional encoding with an MLP that maps coordinates to embeddings results in poor performance ((x,y,z)+MLP), which indicates the amount of high-frequency spatial information in fMRI signals.
@@ 4.7 Visualizations and Interpretations
We found that some queries primarily focus on processing single brain regions, such as Parahippocampal Place Area (PPA) (Figure 6(a)) and Fusiform Face Area (FFA) (6(b)). As previous research has shown, PPA is related to conceptual association, semantic processing and environmental memory (Epstein et al., 1999; Köhler et al., 2002; Bar et al., 2008; Epstein & Ward, 2010) and FFA is known for its critical role in expertise recognition, social cognition and identity memory (Schultz et al., 2003; Tsantani et al., 2021; Xu, 2005). Both are important brain regions for the conceptualization of visual information and are responsible for the interaction between real-time stimulus and past memory (Brewer et al., 1998; Ranganath et al., 2004; Golarai et al., 2007).
Moreover, there are some queries that attend to multiple brain regions, revealing the information transmission between low- and high-level brain regions. For instance, interactions between early visual areas and higher-level regions like PPA and IntraParietal Sulcus (IPS) (Figure 6(e)), revealing a potential pattern for human attention-guided actions (Tunik et al., 2007; Connolly et al., 2016). Additionally, queries are also found responsible for communications between high-level brain regions (Figure 6(c),6(f)). Together, these findings indicate that the learnable queries may reflect the dynamics of human brain activities in the visual task, from seeing and thinking about the image to pressing the button for the visual recall task in NSD (Allen et al., 2022).
We also provide qualitative analysis of model responses in Appendix D.
@@ A.1 Details of each dataset in brain instruction tuning
| Visual Genome QA (Krishna et al., 2017) | Answer image-based questions which require richer semantic understanding of the image than COCO-QA. |
@@
| OK-VQA (Marino et al., 2019) | Answer image-based questions that requires external knowledge beyond the image itself. |
@@
| ST-VQA (Biten et al., 2019) | Answer questions of high-level semantic information present in images as textual cue |
@@
| TDIUC (only used in downstream) (Kafle & Kanan, 2017) | Generate answers for image-based questions from 12 different task directed question types such as object presence and utility/affordance. |
@@ A.3 Instructions
Table 7 lists instructions for each dataset (i.e. ). Some instructions are inspired by (Han et al., 2024b). For question-answering tasks, the instruction is appended after the question in a new line.
@@ Appendix B Implementation details
We choose Vicuna-7b (Zheng et al., 2023) as our backbone LLM. During the brain instruction tuning stage, We use AdamW as the optimizer, with the learning rate of , weight decay of and , . We do not use a learning rate scheduler. We set the batch size to . The instruction tuning is conducted on a machine with L40S GPUs. And each downstream fine-tuning is conducted on a single L40S GPU with a learning rate and batch size. For generations, we have adopted the greedy decoding strategy.
The hyperparameters of the model architecture can be found in Table 8.
| Notation | meaning | value |
| number of fMRI tokens | ||
| / | number of MLP layers in the fMRI encoder | |
| number of queries | ||
| dimension of the query embeddings |
Appendix C Computation Complexity
In the neuroscience-informed attention, the complexity of the dot product between queries and keys is . The complexity of the aggregation of values is , which is neglectable. The MLP maps the hidden representation of dimension to , therefore its complexity is . Therefore, the complexity of the fMRI encoder is given that .

### v2

@@ Abstract
Decoding functional magnetic resonance imaging (fMRI) signals into text has been a key challenge in the neuroscience community, with the potential to advance brain-computer interfaces and uncover deeper insights into brain mechanisms. However, existing approaches often struggle with suboptimal predictive performance, limited task variety, and poor generalization across subjects. In response to this, we propose MindLLM, a model designed for subject-agnostic and versatile fMRI-to-text decoding. MindLLM consists of an fMRI encoder and an off-the-shelf LLM. The fMRI encoder employs a neuroscience-informed attention mechanism, which is capable of accommodating subjects with varying input shapes and thus achieves high-performance subject-agnostic decoding. Moreover, we introduce Brain Instruction Tuning (BIT), a novel approach that enhances the model’s ability to capture diverse semantic representations from fMRI signals, facilitating more versatile decoding. We evaluate MindLLM on comprehensive fMRI-to-text benchmarks. Results demonstrate that our model outperforms the baselines, improving downstream tasks by , unseen subject generalization by , and novel task adaptation by . Furthermore, the attention patterns in MindLLM provide interpretable insights into its decision-making process. Code is available at https://github.com/Graph-and-Geometric-Learning/MindLLM.
@@ 1 Introduction
Another critical challenge lies in designing a unified and subject-agnostic architecture. Current methods of brain multimodal decoding mostly rely on a preprocessing step: selecting responsive voxels by comparing the task-based fMRI to the resting-state fMRI. The selection typically results in higher performance compared to using whole-brain data. However, the varying number and irregular spatial distribution of selected voxels across subjects pose significant challenges for developing a unified architecture. To this end, recent studies (Wang et al., 2024a, b) have proposed pooling or sampling voxels to standardize input dimensions. However, as illustrated in Figure 3, these methods still suffer from the loss of spatial information and uneven representations of certain areas, ultimately compromising performance.
@@ 2 Related Works
This line of research mostly focuses on decoding perceived visual stimuli into natural language from fMRI signals. MindGPT (Chen et al., 2023a), UniBrain(Mai & Zhang, 2023) and BrainCap (Ferrante et al., 2023) employ an fMRI encoder guided by CLIP (Radford et al., 2021) and use a language model (Radford et al., 2019; Wang et al., 2022) to decode natural language from the encoded representations. BrainChat (Huang, 2024) utilizes multiple pretraining strategies (Devlin, 2018; He et al., 2022; Yu et al., 2022) to align fMRI with image and text embeddings. These methods fall short in performance and versatility. UMBRAE (Xia et al., 2024) proposes to learn a mapping from fMRI to stimulus images, which later serves as a proxy input for an off-the-shelf visual language model (VLM). Although they achieve performance improvements, the strategy prevents the model from performing tasks that are not directly related to the stimulus images (e.g., answering memory-related questions). In contrast, our end-to-end Brain Instruction Tuning (BIT) ensures seamless and versatile fMRI-to-text decoding, offering the potential to tackle tasks beyond vision-related ones.
@@
Multimodal fMRI Decoding
Beyond text, recent studies have explored decoding brain signals into other modalities, including images (Scotti et al., 2024b; Wang et al., 2024a; Chen et al., 2023c), videos (Sun et al., 2024; Chen et al., 2023d), speech (Chen et al., 2024) and music (Denk et al., 2023), as well as in reverse directions (Toneva & Wehbe, 2019). The encoder architectures in these works typically either adopt transformers for time-series data, which is distinct from our voxel-based setting, or fall into one of the architectures we have discussed in previous paragraphs. Therefore, focusing on text generation, our work improves the fMRI encoder architecture for voxel-level inputs, and designs novel objectives (i.e., brain instruction tuning) for our target modality.
@@ 3.1 Method Overview
As mentioned before, currently most models for fMRI decoding can not handle varying input shapes and are not subject-agnostic, with only a few exceptions (Wang et al., 2024b). However, these exceptions still suffer from information loss and uneven representations of certain brain areas. To this end, we propose a novel neuroscience-informed attention mechanism to accommodate varying voxel numbers across subjects, enabling a subject-agnostic encoding strategy. Below we talk about the design of queries , keys and values in the attention layer. For values, we directly use the fMRI signal of each voxel, which means . Making each voxel a value token maximally prevents information loss compared to pooling- (Wang et al., 2024a) or sampling-based (Mai & Zhang, 2023) methods. The queries are randomly initialized and learnable. We expect each query to represent a certain pattern of the brain (refer to visualizations in Section 4.8). The design of keys will be discussed below.
Exclude fMRI values from keys The vanilla cross attention (Zhu et al., 2020; Vaswani, 2017) derives both keys and values from the same input source. However, we found this would lead to poor performance in fMRI. We argue the reason: different from images or text, which are usually considered translation-invariant, the positions of voxels carry specific brain functional information, as voxels in different areas are associated with distinct brain functions. Consequently, a voxel’s position alone can theoretically serve as effective keys for attention weight computation (McManus, 2002; Zhang et al., 2023c; Liu et al., 2024). Including fMRI values into keys, however, introduces additional noise instead of valuable information, thus resulting in poorer performance. Moreover, since brain regions tend to serve similar functions across individuals, decoupling voxel positions from fMRI signals can facilitate the sharing of priors across subjects, potentially improving generalization to unseen subjects.
@@
Combining Multiple Parcellations It is crucial to choose an appropriate brain region parcellation. Previous region-based methods (Qiu et al., 2023; Li et al., 2021; Kan et al., 2022) can usually only choose one arbitrarily. In contrast, our model design allows us to combine multiple parcellations by concatenating their respective region encodings to the key embeddings. In conclusion, the final key embeddings are the concatenation of the positional encoding and multiple region encodings,
@@
In addition, we provide a side-by-side comparison between our encoder and existing ones in Appendix F.
@@ 3.3 Brain Instruction Tuning (BIT)
| Method | subject agnostic | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDEr | SPICE |
| SDRecon (Takagi & Nishimoto, 2023) | ✗ | ||||||||
| OneLLM (Han et al., 2024a) | ✗ | ||||||||
| UniBrain (Mai & Zhang, 2023) | ✗ | ||||||||
| BrainCap (Ferrante et al., 2023) | ✗ | ||||||||
| BrainChat (Huang, 2024) | ✗ | ||||||||
| UMBRAE (Xia et al., 2024) | ✗ | ||||||||
| UniBrain (Wang et al., 2024b) | ✓ | ||||||||
| MindLLM (Ours) | ✓ |
@@ 4.1 Settings
Downstream Datasets The downstream dataset will be discussed within each experiment section. See examples and a short description for all datasets we will use in Appendix A. Implementation details could be found in Appendix B.
@@ 4.3 Versatile Decoding
Baselines Besides baselines that could be adapted to this experiment from the previous one, we further consider the following subject-agnostic models as baselines. 1) MindBridge (Wang et al., 2024a) flattens the voxels and adaptively adjusts the padding and stride to pool the voxels into a fixed dimension. The original implementation of MindBridge has subject-specific parameters. However, since those parameters are of the same size, we make them shared across subjects and thus make the model subject-agnostic. 2) UniBrain (Wang et al., 2024b) samples voxels into a fixed number of groups and employs a transformer where groups are treated as tokens. This UniBrain is unrelated to the UniBrain in the previous section; they just share the same name.
@@
The results on subject 1 are shown in Table 2. Our model outperforms baselines, with an average improvement of . Further, by comparing instruction tuning and from-scratch models, we find that instruction tuning has a significant positive effect, with an average improvement of . The results remain stable across different random seeds; for instance, according to our observations, the BLEU-1 score for paragraph captioning exhibits a maximum of variance. We report experiments on other subjects in Appendix G.
@@ 4.4 Unseen Subject Generalization
Our neuroscience-informed, subject-agnostic design enhances generalization to novel subjects, a crucial factor in real-world applications where training a model for each individual is impractical. To evaluate it, we perform instruction tuning on out of the subjects in the natural scene dataset (Allen et al., 2022), and evaluate generalization on the held-out subject. Table 3 shows our model outperforms two other subject-agnostic baselines in most cases, with an average improvement of compared to the second-best model. We anticipate further performance gains through the application of domain adaptation techniques (Xiao et al., 2023; Ganin & Lempitsky, 2015; Ganin et al., 2016; Gong et al., 2012), which we leave as a direction for future work.
@@ 4.5 Adapting to New Tasks
4.6 Performance scale with the number of subjects
We conducted experiments to evaluate how model performance scales with the number of subjects and report performances on the COCO caption task. We examined both the in-distribution (seen subjects) setting in Figure 5 and the out-of-distribution (held-out subjects) setting in Figure 6. Our results show significant performance improvements as the number of training subjects increases, demonstrating that the model benefits from exposure to more subjects during pre-training.
| COCO Caption | ||||||||
| number of subjects | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDEr | SPICE |
| 1 | ||||||||
| 3 | ||||||||
| 5 | ||||||||
| 7 | ||||||||
| number of subjects | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDEr | SPICE |
| 1 | ||||||||
| 3 | ||||||||
| 5 | ||||||||
| 7 |
@@ 4.7 Ablation Study
We conduct ablation studies on the design of key embeddings in the neuroscience-informed attention module in Figure 6. The results strongly validate our design. The vanilla cross attention (Pos Enc.+fMRI) leads to poor performance, while removing fMRI values from the key embeddings (Pos Enc.) yields a significant improvement. Replacing positional encoding with region encodings (Reg. Enc.) accelerates convergence in the early stages since it is grounded by neuroscientific principles. However, it is eventually outperformed by Pos Enc. due to the lack of finer-grained information. Combining the positional encoding and region encodings (Pos Enc.+Reg Enc.) achieves the best model design. In addition, replacing positional encoding with an MLP that maps coordinates to embeddings results in poor performance ((x,y,z)+MLP), which indicates the amount of high-frequency spatial information in fMRI signals. Additional ablation results on evaluation metrics are provided in Appendix E.
@@ 4.8 Visualizations and Interpretations
We found that some queries primarily focus on processing single brain regions, such as Parahippocampal Place Area (PPA) (Figure 7a) and Fusiform Face Area (FFA) (7b). As previous research has shown, PPA is related to conceptual association, semantic processing and environmental memory (Epstein et al., 1999; Köhler et al., 2002; Bar et al., 2008; Epstein & Ward, 2010) and FFA is known for its critical role in expertise recognition, social cognition and identity memory (Schultz et al., 2003; Tsantani et al., 2021; Xu, 2005). Both are important brain regions for the conceptualization of visual information and are responsible for the interaction between real-time stimulus and past memory (Brewer et al., 1998; Ranganath et al., 2004; Golarai et al., 2007).
Moreover, there are some queries that attend to multiple brain regions, revealing the information transmission between low- and high-level brain regions. For instance, interactions between early visual areas and higher-level regions like PPA and IntraParietal Sulcus (IPS) (Figure 7e), revealing a potential pattern for human attention-guided actions (Tunik et al., 2007; Connolly et al., 2016). Additionally, queries are also found responsible for communications between high-level brain regions (Figure 7c,7f). Together, these findings indicate that the learnable queries may reflect the dynamics of human brain activities in the visual task, from seeing and thinking about the image to pressing the button for the visual recall task in NSD (Allen et al., 2022).
@@ 5 Conclusion
Acknowledgments
We are grateful for funding and computing resource support from Amazon Research, Snap Research, National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317, Yale AI Engineering Research Seed Grants, Yale Office of the Provost, and Yale Wu Tsai Institute.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Details of each dataset in brain instruction tuning
| Visual Genome QA (Krishna et al., 2017) | Answer image-based questions that require richer semantic understanding of the image than COCO-QA. |
@@
| OK-VQA (Marino et al., 2019) | Answer image-based questions that require external knowledge beyond the image itself. |
@@
| ST-VQA (Biten et al., 2019) | Answer questions of high-level semantic information present in images as the textual cue |
@@
| TDIUC (only used in downstream) (Kafle & Kanan, 2017) | Generate answers for image-based questions from 12 different task-directed question types such as object presence and utility/affordance. |
@@ A.3 Instructions
Table 9 lists instructions for each dataset (i.e. ). Some instructions are inspired by (Han et al., 2024b). For question-answering tasks, the instructions are appended after the question in a new line.
@@ Appendix B Implementation details
We choose Vicuna-7b (Zheng et al., 2023) as our backbone LLM. During the brain instruction tuning stage, We use AdamW as the optimizer, with the learning rate of , weight decay of and , . We do not use a learning rate scheduler. We set the batch size to . The instruction tuning is conducted on a machine with L40S GPUs for days. And each downstream fine-tuning is conducted on a single L40S GPU with a learning rate and batch size. For generations, we have adopted the greedy decoding strategy.
We use as the number of fMRI tokens. We use a -layer MLP in the fMRI encoder. We set the number of queries to . The dimension of the query embeddings is .
Appendix C Computational Complexity and Runtime Analysis
In the neuroscience-informed attention, the complexity of the dot product between queries and keys is . The complexity of the aggregation of values is , which is negligible. The MLP maps the hidden representation of dimension to , therefore its complexity is . Therefore, the complexity of the fMRI encoder is given that .
Table 10 summarizes the computational complexity and runtime comparison of different models. Despite its superior performance, our model introduces only a marginal increase in encoder-side computational cost compared to other baselines. Notably, the majority of inference time and complexity arises from the LLM component, which is shared across all models. This highlights that the design choices in the encoder, while crucial for performance, do not significantly affect runtime.
| Module | Forward Pass (ms/sample) | Complexity | Notation Explanation | Approximate Values | ||||||||
| UMBRAE |
| MindBridge |
| UniBrain |
| MindLLM |
| LLM | - | - | - |
@@ Appendix D Qualitative Analysis
Appendix E Ablation Studies on Downstream Tasks
Table 11 reports the results of the ablation study on each downstream task. The results validate the design of neuroscience-informed attention.
| + MLP | Pos Enc. + fMRI | Pos Enc. | Reg Enc. | Pos Enc. + Reg Enc. | ||
| COCO-QA | Accuracy | |||||
| VG-QA | Accuracy | |||||
| VQA-v2 | Accuracy | |||||
| A-OKVQA | Accuracy | |||||
| ST-VQA | ANLS | |||||
| Tally QA | Accuracy | |||||
| RMSE | ||||||
| COCO-Caption | BLEU-1 | |||||
| BLEU-2 | ||||||
| BLEU-3 | ||||||
| BLEU-4 | ||||||
| METEOR | ||||||
| ROUGE | ||||||
| CIDEr | ||||||
| SPICE | ||||||
| Paragraph Caption | BLEU-1 | |||||
| BLEU-2 | ||||||
| BLEU-3 | ||||||
| BLEU-4 | ||||||
| METEOR | ||||||
| CIDEr | ||||||
| VQA-E | Accuracy | |||||
| BLEU-1 | ||||||
| BLEU-2 | ||||||
| BLEU-3 | ||||||
| BLEU-4 | ||||||
| CIDEr | ||||||
| METEOR | ||||||
| ROUGE | ||||||
| Average | Perplexity |
Appendix F Architecture Comparison
We make a side-by-side comparison between our model and related baselines in Table 12. Our model is the only one that achieves the four properties listed in the table’s head.
| Method | cross-subject | unified architecture | shared parameters | using all voxels |
| MindEye (Scotti et al., 2024a) | ✗ | ✗ | ✗ | ✓ |
| MindEye2 (Scotti et al., 2024b) | ✓ | ✗ | ✗ | ✓ |
| UMBRAE (Xia et al., 2024) | ✓ | ✗ | ✗ | ✓ |
| MindBridge (Wang et al., 2024a) | ✓ | ✓ | ✗ | ✗ |
| UniBrain (Wang et al., 2024b) | ✓ | ✓ | ✓ | ✗ |
| MindLLM | ✓ | ✓ | ✓ | ✓ |
Appendix G Experiments on Other Subjects
We report experiments on subjects 2, 5, 7 in Table 13. We chose these subjects because they are commonly used in previous works. We observe consistent performance gains over the baselines in the majority of scenarios.
| subject 2 | subject 5 | subject 7 | |||||||||||
| UMBRAE | UniBrain | MindBridge | MindLLM | UMBRAE | UniBrain | MindBridge | MindLLM | UMBRAE | UniBrain | MindBridge | MindLLM | ||
| COCO-QA | Accuracy | ||||||||||||
| VG-QA | Accuracy | ||||||||||||
| VQA-v2 | Accuracy | ||||||||||||
| A-OKVQA | Accuracy | ||||||||||||
| ST-VQA | ANLS | ||||||||||||
| OK-VQA | Accuracy | ||||||||||||
| TallyQA | Accuracy | ||||||||||||
| RMSE | |||||||||||||
| Paragraph Caption | BLEU-1 | ||||||||||||
| BLEU-2 | |||||||||||||
| BLEU-3 | |||||||||||||
| BLEU-4 | |||||||||||||
| METEOR | |||||||||||||
| CIDEr | |||||||||||||
| VQA-E | Accuracy | ||||||||||||
| BLEU-1 | |||||||||||||
| BLEU-2 | |||||||||||||
| BLEU-3 | |||||||||||||
| BLEU-4 | 7.99 | ||||||||||||
| CIDEr | |||||||||||||
| METEOR | |||||||||||||
| ROUGE | |||||||||||||
| FSVQA | VQA Acc. | ||||||||||||
| FSVQA Acc. | |||||||||||||
| BLEU-1 | |||||||||||||
| BLEU-2 | |||||||||||||
| BLEU-3 | |||||||||||||
| BLEU-4 | |||||||||||||
| METEOR | |||||||||||||
| CIDEr | |||||||||||||
Appendix H Latent Space Visualization
To assess how subject-specific and subject-agnostic information evolves through the model, we visualize the latent embeddings at various stages of the encoder in Figure 10. The original inputs exhibit clear subject-specific clusters. After passing through the neuroscience-informed attention layer, the embeddings remain separable by subject, indicating that subject identity is still preserved. As the data flows through successive layers of the MLP, the subject-specific patterns become increasingly mixed, ultimately forming a subject-agnostic representation. This transformation facilitates generalization on held-out subjects in a zero-shot manner.
Appendix I Brain Attention Maps of Other Subjects
We visualize the flat brain attention maps (corresponding to Figure 7a) in Figure 11. We observe that the attention maps exhibit highly similar spatial patterns across subjects, indicating that the model captures shared structural-functional correspondences across human brains, which is expected due to the overall anatomical and functional similarity among individuals. At the same time, we do observe moderate subject-specific variations in the attention maps. These reflect differences in voxel-level functional signals and individual variability in precise voxel locations. The model is able to accommodate these differences through flexible attention mechanisms.